A high-speed drill reaches 2000 rpm in 0.54 s Part A What is the drills angular acceleration? Express your answer to two significant figures and include the appropriate units aValue Units Submit Request Answer ▼ Part B Through how many revolutions does it turn during this first 0.54 s? Express your answer to two significant figures and include the appropriate units IA 6-1 Value Units Submit Request Answer
Answer #1

Here,

angular speed , w = 2000 rpm

w = 2000 * 2pi/60 = 209.4 rad/s

drill's angular acceleration = angular speed/time taken

drill's angular acceleration = (209.4/.54)

drill's angular acceleration = 388 rad/s^2

part B)

for the angle rotated

angle rotated = 0.50 * a * t^2

angle rotated = 0.50 * 388 * 0.54^2

angle rotated = 56.5 radian

angle rotated = 9

the number of revolutions is 9